OpenStudy (anonymous):

add and simplify, if possible x^2+13x/x^2-5x + x^2-6x/x^2-5x

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x^2+13x}{x^2-5x} + \frac{x^2-6x}{x^2-5x}\] \[\Large \frac{x^2+13x+x^2-6x}{x^2-5x}\] \[\Large \frac{2x^2+7x}{x^2-5x}\] \[\Large \frac{x(2x+7)}{x(x-5)}\] \[\Large \frac{\cancel{x}(2x+7)}{\cancel{x}(x-5)}\] \[\Large \frac{2x+7}{x-5}\] So \[\Large \frac{x^2+13x}{x^2-5x} + \frac{x^2-6x}{x^2-5x}=\frac{2x+7}{x-5}\]